What is the viscous drag acting on a sphere with a diameter of 16.2 mm falling through water at 2.3 ms-1? Take the viscosity of the water to be 1.1 x 10-3 Nsm-2.

A 3.9 x 10-4 N

B 7.7 x 10-4 N

C 0.39 N

D 0.77 N


1
Expert's answer
2022-03-24T11:00:41-0400

The viscous drag acting on a sphere is given by

"F_d=6\pi\eta Rv"

"F_d=6*3.14*1.1*10^{-3}*8.1*10^{-3}*2.3\\
=3.9*10^{-4}\:\rm N"

Answer: A